The pressure of 3 moles of an ideal gas at a temperature of 27 having a volume of
10-3m3
1
Expert's answer
2020-06-02T17:54:38-0400

The equation of the ideal gas is:

"pV = \mu RT"

where "p" is the pressure, "V" is the volume, "\mu" is the number of moles, "T" is the absolte temperature and "R = 8.31 \dfrac{J}{mol\cdot K}" is the ideal gas constant.

Expressing the pressure from this equation get:


"p = \frac{\mu RT}{V} = \frac{3\cdot 8.31\cdot (27 + 273)}{10^{-3}} = 7.479\cdot 10^{6} Pa = 7.479 MPa"

Answer. p = 7.479 MPa.